OpenStudy (anonymous):

type the expression as radicals s^6/5

OpenStudy (anonymous):

\[s ^{\frac{ 6 }{ 5 }}=\left( s ^{6} \right)^{\frac{ 1 }{ 5 }}=\sqrt[5]{s ^{6}}\]

OpenStudy (anonymous):

thank you so much! youre awsome

OpenStudy (anonymous):

you are welcome.